OpenStudy (anonymous):

The probability of event A is x, and the probability of event B is y. If the two events are independent, which of these conditions must be true?

OpenStudy (anonymous):

A. P(B|A) = y B. P(A|B) = y C. P(B|A) = x D. P(A and B) = x + y E. P(A and B) = x P(A) y

OpenStudy (kropot72):

If two events A and B are independent, then the probability of A occurring is unaffected by whether or not B has occurred. Therefore: P(A|B) = P(A) Can you now see which conditional probability choice is correct?